OpenStudy (anonymous):

Terry can paint a house in 45 hours. His father can paint the same house in 50 hours and his brother can paint it in 40 hours. How long will it take all three working together to do the painting? Help please? I need the solution. THANKS!

OpenStudy (anonymous):

gimmick is to put the total job as 1 and and each rate as \[\frac{1}{45}+\frac{1}{50}+\frac{1}{40}\]and solve \[(\frac{1}{45}+\frac{1}{50}+\frac{1}{40})t=1\] for t

OpenStudy (anonymous):

t/45 + t/50 + t/40 = 1 121t/1800=1 t=14.88 hrs

OpenStudy (anonymous):

yap. :)

OpenStudy (anonymous):

maybe multiply through by 1800 to get \[40t+36t+45t=1800\] \[121t=1800\] \[t=\frac{1800}{121}\]

OpenStudy (anonymous):

thanks again!